One model airplane travels 30 kilometers per hour faster than another. The faster plane can travel 400 kilometers in the same amount of time the slower plane travels 250 kilometers. Find the speed of both planes.
Answer by Paul(988)

Slow airplanes' speed = x
Fast airplane's speed = x+30
equation:
250(x+30)=400(x)
250x+750=400x
150x=750
x=5
5+30=35
Hence, speed of the slow palne is 5mph and speed of the faster plane is 35mph.
Paul.